OpenStudy (anonymous):

Raul has to choose 3 of the 5 toppings for his frozen yogurt. How many different combinations of toppings are possible? 10 20 60 120

OpenStudy (anonymous):

10 5 choose 3 = 120 / (6 * 2) = 120 / 12 = 10

OpenStudy (anonymous):

Ohhhhh Okay thanks(: